Provide examples (using simple formulations with couple of constraints) of linear programming formulations for

1. Infeasible solution

2. Unbound solution
Answer #1

1.

An example of infeasible solutions

Max 5x + 2y

ST

2x + 3y >= 30

3x + 3y <= 20

x, y >= 0

Here the areas of the inequalities do not overlap in the positive quadrant. As a result it is infeasible

2.

An example of unbound solution

Max 5x + 2y

ST

2x + 3y >= 30

3x + 3y >= 20

x, y >= 0

Here, the inequalities are on the greater side and the objective function is also maximization. Thus the solution is unbound. It can be increased to indefinitely.